simplify the expression: sin5pie/3 cos5pie/3-cos5pie/3 sin pie/3
Found 2 solutions by mamiya, Alan3354:
Answer by mamiya(56)

you say sin(5pie/3) cos(5pie/3)-cos(5pie/3) sin (pie/3), but i think the question was rather
sin(5pie/3) cos(pie/3)-cos(5pie/3) sin (pie/3)
And if the question was what i just wrote , you do it this way.
first let take a look a those trigonometric formulas
sin(a+b) = sin(a) cos(b) + sin(b) cos(a)
sin(a-b) = sin(a) cos(b) - sin(b) cos(a)
Cos(a+b) = cos(a) cos(b) - sin(a) sin(b)
Cos(a-b) = cos(a) cos(b) + sin(a) sin(b)

Now, let's get back to our problem
let's call 5pie/3 a and call pie/3 b
Then, sin5pie/3 cospie/3-cos5pie/3 sin pie/3= sin (a) cos(b)- cos(a) sin(b),
This looks like the second formula up above
so we get sin5pie/3 cospie/3-cos5pie/3 sin pie/3 = Sin ( a-b)
and we know for this case that a = 5pie\3 and b=pie/3
so, sin5pie/3 cospie/3-cos5pie/3 sin pie/3 = Sin ( 5pie/3 - pie/3)
= sin ( 4pie/3)
= -(sqrt(3))/2
So the anwer is Sin( 4pie/3) and if you want to go further and find its value it is -(sqrt(3))/2


Answer by Alan3354(69443)

simplify the expression: sin5pie/3 cos5pie/3-cos5pie/3 sin pie/3
----------
It's the Greek letter pi, not the dessert, pie.